iii. Suppose that the price of one ear of corn was $0.05 in 1920, that the CPI in 1920 was 10, and that in 1990 the CPI was 180. What is the price of a 1920 ear of corn in 1990 dollars?
iiii. In 1954, Mickey Mantle earned $21,000 playing for the New York Yankees. The CPI in 1954 was 26.9, and the CPI in 2010 was 218.06. What is Mickey Mantle’s 1954 salary in 2010 dollars?
1
Expert's answer
2020-07-22T10:15:10-0400

"\forall" $ 0.05 =10 CPI in 1920

"\therefore" how many $ =180 CPI in 1990.

0.05=10

?=180

0.05×180=9

9 ÷10=0.9

"\therefore" ,the price of a 1920 ear of corn in 1990 dollars is $ 0.90


(ii)26.9 CPI = $21,000

218.06 CPI= $ ?

"\therefore" , $21,000×218.06=4,579,260

4,579,260÷26.9=170,232.71.

Mickey Mantle's 1954 salary in 2010 dollars is $ 170,232.71